The following data show the brand, price ($), and the overall score for 6 stereo headphones that were tested by Consumer Reports. The overall score is based on sound quality and effectiveness of ambient noise reduction. Scores range from 0 (lowest) to 100 (highest). The estimated regression equation for these data is = 24.881 + 0.301 x, where x = price ($) and y = overall score.

Brand Price ($) Score
Bose 180 75
Scullcandy 160 72
Koss 95 61
Phillips/O'Neill 70 56
Denon 60 40
JVC 35 26

Round your answers to three decimal places.

a. Compute SST, SSR, and SSE.

SST
SSR
SSE

b. Compute the coefficient of determination r 2. Comment on the goodness of fit.

r 2 =

SelectThe least squares line provided a very good fitThe least squares line provided a bad fit

c. What is the value of the sample correlation coefficient?

r =
Answer #1

using Minitab >stat>basic stat>Regression

Regression Analysis: Score versus Price ($)

The regression equation is
Score = 24.88 + 0.3012 Price ($)


S = 8.25343 R-Sq = 84.8% R-Sq(adj) = 81.0%


Analysis of Variance

Source DF SS MS F P
Regression 1 1519.52 1519.52 22.31 0.009
Error 4 272.48 68.12
Total 5 1792.00

a. Compute SST, SSR, and SSE.

SST 1792
SSR 1519.52
SSE 272.48

b. the coefficient of determination r 2

r 2 = 84.8%

the least squares line provided a very good fit

c. the sample correlation coefficient is

r =0.921